# Morphology of 32 Repeating Fast Radio Burst Sources at Microsecond Time Scales with CHIME/FRB

### v1

@@ Abstract
The Canadian Hydrogen Intensity Mapping Experiment Fast Radio Burst (CHIME/FRB) project has discovered the most repeating fast radio burst (FRB) sources of any telescope. However, most of the physical conclusions derived from this sample are based on data with a time resolution of 1 ms. In this work, we present for the first time a morphological analysis of the raw voltage data for 118 bursts from 32 of CHIME/FRB’s repeating sources. We do not find any significant correlations amongst fluence, dispersion measure (DM), burst rate, and burst duration. Performing the first large-scale morphological comparison at timescales down to microseconds between our repeating sources and 125 non-repeating FRBs, we find that repeaters are narrower in frequency and broader in duration than non-repeaters, supporting previous findings. However, we find that the duration-normalized sub-burst widths of the two populations are consistent, possibly suggesting a shared physical emission mechanism. Additionally, we find that the spectral fluences of the two are consistent. When combined with the larger bandwidths and previously found larger DMs of non-repeaters, this suggests that non-repeaters may have higher intrinsic specific energies than repeating FRBs. We do not find any consistent increase or decrease in the DM ( pc cm-3 yr-1) and scattering timescales ( ms yr-1) of our sources over year periods.
@@ 1 Introduction
Fast radio bursts (FRBs) are brief () bursts of radio emission originating from primarily extragalactic distances with radio luminosities erg s-1 (Lorimer et al., 2007). FRBs are typically divided into non-repeating (or “one-off”) FRBs (CHIME/FRB Collaboration et al., 2021) and repeating FRBs (Spitler et al., 2016; CHIME/FRB Collaboration et al., 2019a, b, 2023a; Fonseca et al., 2020). While most FRBs’ repeatability is seemingly uncorrelated, one source shows definitive periodic windows of burst activity (CHIME/FRB Collaboration et al., 2020a).
Despite the large number of published FRBs, the progenitors and emission mechanisms of these sources remain uncertain. The detection of an FRB-like burst from the Galactic magnetar SGR 1935+2154 lent strong evidence to magnetars as possible progenitors (CHIME/FRB Collaboration et al., 2020b; Bochenek et al., 2020). However, the discovery of a repeating FRB in the globular cluster of M81 challenged this hypothesis, as magnetars are unexpected in such old environments (Bhardwaj et al., 2021; Kirsten et al., 2022).
Theoretical models for the emission mechanism generally involve neutron stars (NS), and fall into two main categories: those that suggest emission arises from a synchrotron maser process driven by shocks far from the central compact object (Metzger et al., 2019), and those that propose the emission originates closer to the central compact object, such as within the magnetosphere (Lu & Kumar, 2018; Ioka, 2020). Recent evidence has increasingly supported the idea of a magnetospheric origin for at least some sources (CHIME/FRB Collaboration et al., 2020a; Nimmo et al., 2021, 2022; Mckinven et al., 2024; Nimmo et al., 2024).
@@
Detect s or ns-scale features: Some FRBs exhibit extremely short timescale features known as micro features and nano-shots (Farah et al., 2018; Cho et al., 2020; Nimmo et al., 2021; Majid et al., 2021; Nimmo et al., 2022; Hewitt et al., 2023). Such features again strongly support magnetospheric emission models for at least for the sources that exhibit them (Nimmo et al., 2021). There has not, however, been a large-scale analysis of the microstructure of repeating versus non-repeating sources.
@@
In this paper, we present the first large-scale study of the morphology of repeating FRBs at time resolutions down to 2.56 using the CHIME/FRB experiment. In Section 2, we discuss the CHIME/FRB instrument, its complex voltage recording system, and the analysis pipeline used to determine burst properties. In Section 3, we present the sources analyzed in this work. In Section 4, we present various analyses related to the burst properties. We start by searching for correlations amongst various morphological parameters for our repeating FRB dataset. We then compare the morphological parameters of the repeating FRBs with those of the non-repeating FRBs studied by Sand et al. (2024). Finally, we explore whether any of our repeaters show changing environments with time. In Section 5, we discuss our results in the context of the FRB population. We end in Section 6 by summarizing our work.
@@ 2.2 Analysis Pipeline
Step 2: The data are downsampled from a time resolution of 2.56 such that a minimum peak S/N of is reached for each burst. The S/N limit is chosen on a burst-to-burst basis given burst characteristics. If a burst has not reached the desired S/N threshold at a time resolution of 0.655 ms, we either a) remove the burst from our sample or b) include the burst but caution that the parameter fit (discussed below) may be less reliable.
Step 3: We determine a DM by maximizing S/N for the burst. We use this DM as an input to search for a structure-optimized DM using the DM phase routine (Seymour et al., 2019).
Step 4: We determine a profile for the burst by summing over all non-RFI channels. We then smooth the profile using a locally weighted scatter plot smoothing method (LOWESS) implemented by statsmodel in python (Seabold & Perktold, 2010). We then determine the number of peaks within the burst using the find_peaks algorithm implemented by SciPy (Virtanen et al., 2020). In certain cases, we manually supply the number of peaks when the find_peaks algorithm does not pick out the same number of peaks as determined by eye.
@@ 2.3 Burst Rates
We follow the methodology outlined by CHIME/FRB Collaboration et al. (2023a) to determine burst rates. In brief, we calculate the total exposure time for each source based on its improved position (CHIME/FRB Collaboration et al., 2023a; Michilli et al., 2023), and determine the number of bursts found within the FWHM of their detection beam at 600 MHz. We assume a Poissonian uncertainty on the burst rate, as this dominates over the uncertainty in exposure due to the improved source positions from the beam-formed data.
We calculate a spectral fluence threshold per source using the method outlined by Josephy et al. (2019). The fluence threshold accounts for three factors - day-to-day instrument gain variations, changes in synthesized beam response during source transit, and varying emission bandwidths and frequency centers. We simulate thresholds for different detection scenarios to estimate relative sensitivity between simulated and real detections, scaling the initial fluence threshold inferred from the real detection by the simulated thresholds. For repeaters, each realization of the simulation uses a randomly selected burst as the reference for determining relative sensitivity and initial threshold. Using the 95th percentile of the scaled threshold distribution per source, we scale each source’s burst rate to a fluence threshold of 5 Jy ms assuming a power-law index of for the cumulative energy distribution.
@@ 2.4 Biases
The CHIME/FRB system for recording the complex raw voltages has a number of biases in addition to those discussed for the real-time system by Merryfield et al. (2023). The full data can only be saved for sources with a DM pc cm-3 due to the size of the buffer for this system, with channels at the top of the band lost. Three sources in our sample, FRB 20200809E, FRB 20190417A, and FRB 20190216A, have a DM pc cm-3. Second, only ms of data are saved around a given event (after de-dispersion). Thus, events significantly longer than 100 ms, or events for which the initial DM was inaccurately determined, can be clipped. Certain events also lack data at all frequency channels due to issues with the early recording system. Additionally, only events above a given S/N threshold will trigger this recording system. This threshold is different for known-repeating (S/N ) and non-repeating FRBs (S/N ). In both cases, faint bursts or bursts with particularly complex structure for which the peak S/N is low may be missed. While some of the selection biases of CHIME/FRB have been described by Merryfield et al. (2023), the selection biases against complex, narrow-band bursts (such as those studied here) has not been well studied. Lastly, Sand et al. (2024) found that bursts studied at time resolutions ms tended to have higher peak flux densities. This suggested that CHIME/FRB may be missing ultra-narrow bursts that do not have such high peak flux densities due to the averaging of these bursts with noise at time resolutions of 0.983 ms.
@@ 3 Repeating FRB Sources
The CHIME/FRB collaboration previously published 45 repeating sources of FRBs (CHIME/FRB Collaboration et al., 2019a, b, 2023a; Fonseca et al., 2020; Bhardwaj et al., 2021; Lanman et al., 2022). Here, we focus on the sources published in the above works that have at least one CHIME-detected burst with the complex raw voltages recorded. We do not include FRB 20180916B in our work as there has already been significant analysis of this FRB at high-time resolution with CHIME/FRB (Sand et al., 2023). We exclude two highly active, previously published repeaters detected by CHIME/FRB: FRB 20201124A and FRB 20220912A. We also do not include bursts for which the S/N at a time resolution of 0.655 ms is . The remainder amounts to 118 bursts with recorded complex raw voltages from 32 repeating FRBs between 2018 August and 2023 September. The end date is chosen to be consistent with the cutoff for the second CHIME/FRB catalog that is in preparation.
We present a sub-sample of the frequency versus time (hereafter referred to as ‘waterfall’) plots in the Appendix in Figure A1. We also list each burst and its fit parameters in Table A1. We could not determine a flux/fluence measurement for four repeater bursts in our sample (FRB 20190626A, FRB 20191217A, FRB 20210302E and FRB 20220203A) as they were detected from the meridian of CHIME/FRB where our primary beam model is not sufficiently well understood. We also could not determine a flux/fluence for FRB 20210130E due to its low S/N. In Table A2, we present the updated burst rates for our sources.
@@ 4.1 Burst Property Correlations
In Figure 1, we plot various combinations of burst rate, average source duration, average source fluence, and extragalactic (+halo) DM in order to search for correlations. While we primarily focus on repeating FRBs for these correlations, we include non-repeaters from Sand et al. (2024) in some of the comparisons for which it makes sense to do so (e.g., correlations not including burst rate). For the average burst duration and fluence, the plotted error bars encapsulate the range of values for that source, e.g., the range of burst durations exhibited by that repeater over all of its bursts.
We define the extragalactic DM as the excess after subtracting the NE2001 Galactic component (Cordes & Lazio, 2002) from the DM determined using fitburst. We assume a 20 uncertainty on the NE2001 DM contribution. While some models and constraints on the Milky Way halo predict significant latitude-dependent structure (e.g., Yamasaki & Totani, 2020; Das et al., 2021), the variation between the median DM halo values from most commonly accepted models is of order that expected from the variations along individual lines-of-sight (Cook et al., 2023). Thus, we do not account for the halo contribution to the observed DMs. This has the same effect on the analysis as would subtracting a constant halo value from each line of sight. We refer to this DM as the extragalactic DM (+halo). We do not account for the Galactic scattering contribution, as it is not additive in the same way as the DM. However, all of our scattering values are significantly larger than the predicted Galactic scattering contribution (Cordes & Lazio, 2002). We also only include sources in our analysis for which there are at least two bursts from that source that have the complex raw voltages recorded. This totals 25 different repeating sources.
@@
To account for uncertainties in our measured parameters, we follow the technique presented by CHIME/FRB Collaboration et al. (2023a). We first perform a Monte Carlo (MC) simulation in which we re-determine a given parameter (e.g., average duration of a source) 1000 times. We do this by sampling from the uncertainty region for each burst for that given parameter in order to determine a new average parameter. We calculate an associated Spearman p-value and correlation coefficient for every iteration, and report the median p-value and correlation coefficient from the MC along with the Median Absolute Deviation From the Median (MADFM).
@@ 4.1.1 Burst Rate & Average Duration
However, as seen in Figure 1, many sources emit bursts spanning approximately an order of magnitude in burst duration, with FRB 20181030A emitting bursts over approximately two orders of magnitude in duration. If beaming angle is proportional to burst width, then this implies a significant change in the beaming angle for an individual source. It is also possible that the large range of burst durations seen from a given source obscures an underlying correlation. Therefore we also re-do the correlation study using only the maximum burst width from a given source (see Figure A2 in the Appendix). The median p-value from 1000 MC samples for this correlation is 0.096. While this is smaller than that using average burst duration, it is still not significant.
@@ 4.1.3 Average Duration & Extragalactic DM Correlation
As seen in the bottom left panel of Figure 4.1, we do not find a significant correlation between the average duration of our sources and their extragalactic (+halo) DM, as might be expected if higher-DM sources are on average farther away and undergo greater cosmological time dilation (Connor, 2019). We also re-do this analysis including the non-repeating bursts from Sand et al. (2024) as this analysis does not require a burst repetition rate. The p-value for the correlation using this additional sample of bursts is 0.98, which is not significant (see Figure A3 in the Appendix).
@@ 4.1.4 Average Fluence & Extragalactic DM Correlation
As seen in the lower left panel of Figure 1, we do not find a correlation between the average fluence of our sources and their extragalactic (+halo) DM, as might be expected if higher-DM sources are on average farther away and all repeating FRBs follow a universal luminosity function with the same average luminosity. We re-do the correlation using the median fluence, but again find no correlation. We also re-do the correlation using the non-repeater sample from Sand et al. (2024). The Spearman p-value is 0.005 (See Figure A3 in the Appendix), still larger than our significance threshold of but smaller than that when examining the two samples separately.

### v2

# Morphology of 35 Repeating Fast Radio Burst Sources at Microsecond Time Scales with CHIME/FRB

@@ Abstract
The Canadian Hydrogen Intensity Mapping Experiment Fast Radio Burst (CHIME/FRB) project has discovered the most repeating fast radio burst (FRB) sources of any telescope. However, most of the physical conclusions derived from this sample are based on data with a time resolution of 1 ms. In this work, we present for the first time a morphological analysis of the raw voltage data (Thompson et al., 2017) for 124 bursts from 35 of CHIME/FRB’s repeating sources. We do not find any significant correlations amongst fluence, dispersion measure (DM), burst rate, and burst duration. Performing the first large-scale morphological comparison at timescales down to microseconds between our repeating sources and 125 non-repeating FRBs, we find that repeaters are narrower in frequency and broader in duration than non-repeaters, supporting previous findings. However, we find that the duration-normalized sub-burst widths of the two populations are consistent, possibly suggesting a shared physical emission mechanism. Additionally, we find that the spectral fluences of the two are consistent. When combined with the larger bandwidths and previously found larger DMs of non-repeaters, this suggests that non-repeaters may have higher intrinsic specific energies than repeating FRBs. We do not find any consistent increase or decrease in the DM ( pc cm-3 yr-1) and scattering timescales ( ms yr-1) of our sources over year periods.
@@ 1 Introduction
Fast radio bursts (FRBs) are brief () bursts of radio emission originating from primarily extragalactic distances with radio luminosities erg s-1 (Lorimer et al., 2007). FRBs are typically divided into non-repeating (or “one-off”) FRBs (CHIME/FRB Collaboration et al., 2021) and repeating FRBs (Spitler et al., 2016; CHIME/FRB Collaboration et al., 2019a, b, 2023a; Fonseca et al., 2020). While most FRBs’ repeatability is seemingly uncorrelated, one source shows definitive periodic windows of burst activity (CHIME/FRB Collaboration et al., 2020).
Despite the large number of published FRBs, the progenitors and emission mechanisms of these sources remain uncertain. Theoretical models generally involve neutron stars (NS), and fall into two main categories: those that suggest emission arises from a synchrotron maser process driven by shocks far from the central compact object (Metzger et al., 2019), and those that propose the emission originates closer to the central compact object, such as within the magnetosphere (Lu & Kumar, 2018; Ioka, 2020). Recent evidence has increasingly supported the idea of a magnetospheric origin for at least some sources (CHIME/FRB Collaboration et al., 2020; Nimmo et al., 2021, 2022; Mckinven et al., 2024; Nimmo et al., 2024).
@@
Detect s or ns-scale features: Some FRBs exhibit extremely short timescale features known as micro features and nano-shots (Farah et al., 2018; Cho et al., 2020; Nimmo et al., 2021; Majid et al., 2021; Nimmo et al., 2022; Hewitt et al., 2023). Such features again strongly support magnetospheric emission models for at least the sources that exhibit them (Nimmo et al., 2021).
@@
In this paper, we present the first large-scale study of the morphology of repeating FRBs at time resolutions down to 2.56 using the CHIME/FRB experiment. In Section 2, we discuss the CHIME/FRB instrument, its complex voltage recording system, and the analysis pipeline used to determine burst properties. In Section 3, we present the sources analyzed in this work. In Section 4, we present a variety analyses related to the burst properties. We start by searching for correlations amongst relevant morphological parameters for our repeating FRB dataset. We then compare the morphological parameters of the repeating FRBs with those of the non-repeating FRBs studied by Sand et al. (2024). Finally, we explore whether any of our repeaters show changing environments with time. In Section 5, we discuss our results in the context of the FRB population. We end in Section 6 by summarizing our work.
@@ 2.2 Analysis Pipeline
Step 2: The data are downsampled from a time resolution of 2.56 such that a minimum peak S/N of is reached for each burst. The S/N limit is chosen on a burst-to-burst basis given burst characteristics. If a burst has not reached the desired S/N threshold at a time resolution of 0.655 ms, we either a) remove the burst from our sample if it is particularly faint or b) include the burst but caution that the parameter fit (discussed below) may be less reliable.
Step 3: We determine a DM by maximizing S/N for the burst. We use this DM as an input to search for a structure-optimized DM using the DM phase routine (Seymour et al., 2019). We do not fit for the DM of individual sub-bursts, but instead assume one DM value per event.
Step 4: We determine a profile for the burst by summing over all non-RFI channels. We then smooth the profile using a locally weighted scatter plot smoothing method (LOWESS) implemented by statsmodel in python (Seabold & Perktold, 2010). We then determine the number of peaks within the burst (e.g., number of sub-bursts) using the find_peaks algorithm implemented by SciPy (Virtanen et al., 2020). In certain cases, we manually supply the number of peaks when the find_peaks algorithm does not pick out the same number of peaks as determined by eye.
@@ 2.3 Burst Rates
We follow the methodology outlined by CHIME/FRB Collaboration et al. (2023a) to determine burst rates. In brief, we calculate the total exposure time for each source based on its improved position (CHIME/FRB Collaboration et al., 2023a; Michilli et al., 2023), and determine the number of bursts found within the FWHM of their detection beam at 600 MHz. We do not distinguish between periods of high activity or quiescence. We assume a Poissonian uncertainty on the burst rate, as this dominates over the uncertainty in exposure due to the improved source positions from the beam-formed data.
We calculate a spectral fluence threshold per source using the method outlined by Josephy et al. (2019). The fluence threshold accounts for three factors - day-to-day instrument gain variations, changes in synthesized beam response during source transit, and varying emission bandwidths and frequency centers. We simulate thresholds for different detection scenarios to estimate relative sensitivity between simulated and real detections, scaling the initial fluence threshold inferred from the real detection by the simulated thresholds. For repeaters, each realization of the simulation uses a randomly selected burst as the reference for determining relative sensitivity and initial threshold. Using the 95th percentile of the scaled threshold distribution per source, we scale each source’s burst rate to a fluence threshold of 5 Jy ms assuming a power-law index of for the cumulative energy distribution (CHIME/FRB Collaboration et al., 2023a).
@@ 2.4 Biases
The CHIME/FRB system for recording the complex raw voltages has a number of biases in addition to those discussed for the real-time system by Merryfield et al. (2023). The full data can only be saved for sources with a DM pc cm-3 due to the size of the buffer for this system, with channels at the top of the band lost. Three sources in our sample, FRB 20200809E, FRB 20190417A, and FRB 20190216A, have a DM pc cm-3. Second, only ms of data are saved around a given event (after de-dispersion). Thus, events significantly longer than 100 ms, or events for which the initial DM was inaccurately determined by the real-time pipeline, can be clipped. Certain events also lack data at all frequency channels due to issues with the early recording system. Additionally, only events above a given S/N threshold will trigger this recording system. This threshold is different for known-repeating (S/N ) and non-repeating FRBs (S/N ). In both cases, faint bursts or bursts with particularly complex structure for which the peak S/N is low may be missed. While some of the selection biases of CHIME/FRB have been described by Merryfield et al. (2023), the selection biases against complex, narrow-band bursts (such as those studied here) has not been well studied. Lastly, Sand et al. (2024) found that bursts studied at time resolutions ms tended to have higher peak flux densities. This suggested that CHIME/FRB may be missing ultra-narrow bursts that do not have such high peak flux densities due to the averaging of these bursts with noise at time resolutions of 0.983 ms (CHIME/FRB Collaboration et al., 2023a).
@@ 3 Repeating FRB Sources
The CHIME/FRB collaboration previously published 45 repeating sources of FRBs (CHIME/FRB Collaboration et al., 2019a, b, 2023a; Fonseca et al., 2020; Bhardwaj et al., 2021; Lanman et al., 2022). Here, we focus on the sources published in the above works that have at least one CHIME-detected burst with the complex raw voltages recorded. We do not include FRB 20180916B in our work as there has already been significant analysis of this FRB at high-time resolution with CHIME/FRB (Sand et al., 2023). We exclude two highly active, previously published repeaters detected by CHIME/FRB: FRB 20201124A and FRB 20220912A. Including highly active repeaters might bias certain comparisons as it will dominate the distribution shape, washing away the effect of the other sources. Additionally, highly active repeaters may represent a special sub-class of repeaters, which is not the main focus of this work. We also do not include bursts for which the S/N at a time resolution of 0.655 ms is . The remainder amounts to 124 bursts with recorded complex raw voltages from 35 repeating FRBs between 2018 August and 2023 September. The end date is chosen to be consistent with the cutoff for the second CHIME/FRB catalog (CHIME/FRB Collaboration, 2025).
We present a sub-sample of the frequency versus time (hereafter referred to as ‘waterfall’) plots in the Appendix in Figure A1. We also list each burst and its fit parameters in Table A1. We could not determine a flux/fluence measurement for four repeater bursts in our sample (FRB 20190626A, FRB 20191217A, FRB 20210302E and FRB 20220203A) as they were detected from the meridian of CHIME/FRB where our primary beam model is not sufficiently well understood. Thus, for all analysis requiring fluences, we do not include these bursts. We also could not determine a flux/fluence for FRB 20210130E due to its low S/N. In Table A2, we present the updated burst rates for our sources.
@@ 4.1 Burst Property Correlations
In Figure 1, we plot combinations of burst rate, average source duration, average source fluence, and extragalactic (+halo) DM in order to search for correlations. While we primarily focus on repeating FRBs for these correlations, we include non-repeaters from Sand et al. (2024) in some of the comparisons for which it makes sense to do so (e.g., correlations not including burst rate). For the average burst duration and fluence, the plotted error bars encapsulate the range of values for that source, e.g., the range of burst durations exhibited by that repeater over all of its bursts.
We define the extragalactic DM as the excess after subtracting the NE2001 Galactic component (Cordes & Lazio, 2002) from the DM determined using fitburst. We assume a 20 uncertainty on the NE2001 DM contribution. While some models and constraints on the Milky Way halo predict significant latitude-dependent structure (e.g., Yamasaki & Totani, 2020; Das et al., 2021), the variation between the median DM halo values from most commonly accepted models is of order that expected from the variations along individual lines-of-sight (Cook et al., 2023). Thus, we do not account for the halo contribution to the observed DMs. This has the same effect on the analysis as would subtracting a constant halo value from each line of sight. We refer to this DM as the extragalactic DM (+halo). We do not account for the Galactic scattering contribution, as it is not additive in the same way as the DM. However, all of our scattering values are significantly larger than the predicted Galactic scattering contribution (Cordes & Lazio, 2002). For this specific analysis, we only include sources for which there are at least two bursts from that source that have the complex raw voltages recorded. This totals 25 different repeating sources.
@@
To account for uncertainties in our measured parameters, we follow the technique presented by CHIME/FRB Collaboration et al. (2023a). We first perform a Monte Carlo (MC) simulation in which we re-determine a given parameter (e.g., average duration of a source) 1000 times. We do this by sampling from the uncertainty region for each burst for that given parameter in order to determine a new average parameter. We calculate an associated Spearman p-value and correlation coefficient for every iteration, and report the median p-value and correlation coefficient from the MC along with the Median Absolute Deviation From the Median (MADFM). For repeaters for which the burst rate is zero222When calculating the burst rate, we only include bursts that occurred within the FWHM of a given CHIME/FRB beam at 600 MHz. Thus, it is possibly for a given repeater to have no bursts occur within these parameters, and thus have a burst rate of zero. but the new, sampled rate will not be zero, we do not show them in the below plots but do include them when calculating correlation coefficients and p-values.
@@ 4.1.1 Burst Rate & Average Duration
However, as seen in Figure 1, many sources emit bursts spanning approximately an order of magnitude in burst duration, with FRB 20181030A emitting bursts over approximately two orders of magnitude in duration. If beaming angle is proportional to burst width, then this implies a significant change in the beaming angle for an individual source. It is also possible that the large range of burst durations seen from a given source obscures an underlying correlation. Therefore we also re-do the correlation study using only the maximum burst width from a given source (see Figure A2 in the Appendix). The median p-value from 1000 MC samples for this correlation is 0.68, still not significant.
@@ 4.1.3 Average Duration & Extragalactic DM Correlation
As seen in the bottom left panel of Figure 4.1, we do not find a significant correlation between the average burst duration for our sources and their extragalactic (+halo) DM, as might be expected if higher-DM sources are on average farther away and undergo greater cosmological time dilation (Connor, 2019). We also re-do this analysis including the non-repeating bursts from Sand et al. (2024) as this analysis does not require a burst repetition rate. The p-value for the correlation using this additional sample of bursts is 0.97, which is not significant (see Figure A3 in the Appendix). The p-value when just using the non-repeaters is 0.58 (see also Sand et al., 2024).
@@ 4.1.4 Average Fluence & Extragalactic DM Correlation
As seen in the lower right panel of Figure 1, we do not find a correlation between the average fluence of our sources and their extragalactic (+halo) DM, as might be expected if higher-DM sources are on average farther away and all repeating FRBs follow a universal luminosity function with the same average luminosity. We re-do the correlation using the median fluence, but again find no correlation. We also re-do the correlation using the non-repeater sample from Sand et al. (2024). The Spearman p-value is 0.005 (See Figure A4 in the Appendix), still larger than our significance threshold of but smaller than that when examining the two samples separately. The p-value from the non-repeaters alone is 0.02.